Question 7 of 14> Attempt center of its distribution. Mean Median Mode Midrange height of corn stalks in a field of corn weekly record sales of a one- hit wonder pop group salary of male and female soccer players insulation thickness of copper wire within 10 mm age of lifetime achievement award winners religious affiliation of United States citizens Answer Bank about us careers privacy policyte of use contact us help ) d,英圖20
Answer #1

Mid range has got no particular examples. It can be used for normally distributed data(or approximately normal) without outliers as it comes close to the mean. So, the examples where mid-range can be used as measure of central tendency are:

1. Heights of students in a class.

2. Weight of students in a school.

3. Daily temperatures for a month.

Generally mean is used in these cases but the calculation of midrange is easier than the mean as it considers only maximum and minimum value.

Midrange =(maximum value+minimum value)/2

Example: After removing any outliers, let the heights of 10 students(in feet) be: 4.6, 4.8, 5.4, 5.8, 5.9, 6, 5.8, 5.3, 4.9, 4.7

Mean of the data =(4.6+4.8+.....+4.7)/10 =53.2/10 =5.32

Midrange of the data =(6+4.6)/2 =10.6/2 =5.3​​​​​​

We can see that they are close to each other. And midrange is easier to calculate. Hence, midrange can be used for such data which is normal and no outliers.